# REVS: Unlearning Sensitive Information in Language Models via Rank Editing in the Vocabulary Space

### v5

@@ C.5 Head-Projection Baseline
We followed the original implementation with a modified objective to prevent the target from appearing in the top $k=100$ logit distribution elements. For the Llama-3-8B model, we targeted layers $3$-$5$ and Patil layers $15$-$31$. The SSN dataset configuration used a loss break of $1.9\times 10^{-3}$, loss prediction probability coefficient of $3.08\times 10^{-2}$, learning rate of $1.03\times 10^{-2}$, and $10$ gradient steps. The EMAIL used a loss break of $7.6\times 10^{-4}$, loss prediction probability coefficient of $2.58\times 10^{-1}$, learning rate of $9.19\times 10^{-3}$, and $25$ gradient steps. For the GPT-J-6B model we targeted layers $3$-$8$ and Patil layers $15$-$27$. For the SSN dataset we used a loss break of $6.596\times 10^{-2}$, loss prediction probability coefficient of $3.271$, learning rate of $2.045\times 10^{-1}$, and $25$ gradient steps. The EMAIL dataset configuration used similar layer targeting but modified the hyperparameters, with a loss break of $6\times 10^{-4}$, loss prediction probability coefficient of $6.733\times 10^{-1}$, learning rate of $1.868\times 10^{-1}$, and $5$ gradient steps. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ C.6 Max-Entropy Baseline
For the Llama-3-8B model, we targeted layers $3$-$5$ and Patil layers $15$-$31$. The SSN dataset configuration employed a loss break of $1.99\times 10^{-2}$, loss prediction probability coefficient of $20.5$, learning rate of $7.6\times 10^{-3}$, and $25$ gradient steps. The EMAIL dataset adjusted these parameters, using a loss break of $1.75\times 10^{-3}$, loss prediction probability coefficient of $1.35\times 10^{-2}$, learning rate of $9.2\times 10^{-3}$, and $5$ gradient steps. For the GPT-J-6B model, we targeted layers $3$-$8$ and Patil layers $15$-$27$. For the SSN dataset, we used a loss break of $7.0559\times 10^{-2}$, loss prediction probability coefficient of $25.9$, learning rate of $1.937\times 10^{-1}$, and $5$ gradient steps. The EMAIL dataset used a loss break of $3.634\times 10^{-2}$, loss prediction probability coefficient of $1.907\times 10^{-2}$, learning rate of $6.65\times 10^{-2}$, and $10$ gradient steps. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ D.1 Alternative Neuron Selection Strategies
We explored several alternative methods for neuron selection, besides our hybrid approach that considers both neuron activation and association with the target token (Section 3.1.2). These alternatives include: (a) selecting the most activated neurons given the prompt, (b) selecting neurons solely based on their association with the target token, (c) neurons with highest gradients with respect to the prompt (Dai et al., 2022), and (d) random neuron selection as a baseline. As shown in Table 7, our evaluation on the Emails dataset on GPT-J-6B, using the same seed (00) that was used for finding the best hyper-parameters, demonstrated that the hybrid approach (rank & activation) outperformed these alternatives in achieving superior results in both metrics of unlearning effectiveness and model integrity.
@@ D.2 Alternative Neuron Editing Strategies
We explore an alternative neuron editing strategy inspired by DEPN, where the target neurons are directly zeroed out instead of applying our proposed rank-based editing approach. As shown in Table 8, zeroing neurons results in significantly inferior performance compared to REVS.
@@ E.2 Alternative Token Selection Strategies
We investigated token selection strategies for unlearning on the GPT-J-6B Email dataset. Table 9 shows that selecting the rarest tokens yields the best performance across all metrics. The rarest token approach significantly outperforms other strategies like most frequent, first, or random token selection, demonstrating substantially higher effectiveness in unlearning.
@@ F.2 Synthetic SSN Dataset
We then randomly assigned fake sensitive information to each placeholder. To populate the placeholders, we used Claude 3 Sonnet to generate a list of $40$ unique names (since each sentences will be assigned with the same name and SSN for allowing evaluating generalization). Additionally, we generated $200$ random dates and $40$ random Social Security Numbers (SSNs). To get the model to memorize the sentences containing the sensitive information, we fine-tuned the model with a low learning rate of $62e-$ 6 to minimize unintended impact. The fine-tuning process ran for 4 epochs with a batch size of $1$. Table 12 shows 10 example sentences from this dataset.

### v6

@@ C.5 Head-Projection Baseline
We followed the original implementation with a modified objective to prevent the target from appearing in the top $k=100$ logit distribution elements. For the Llama-3-8B model, we targeted layers 33-55 and Patil layers 1515-$31$. The SSN dataset configuration used a loss break of $1.9\times 10^{-3}$, loss prediction probability coefficient of $3.08\times 10^{-2}$, learning rate of $1.03\times 10^{-2}$, and $10$ gradient steps. The EMAIL used a loss break of $7.6\times 10^{-4}$, loss prediction probability coefficient of $2.58\times 10^{-1}$, learning rate of $9.19\times 10^{-3}$, and $25$ gradient steps. For the GPT-J-6B model we targeted layers 33-88 and Patil layers 1515-$27$. For the SSN dataset we used a loss break of $6.596\times 10^{-2}$, loss prediction probability coefficient of $3.271$, learning rate of $2.045\times 10^{-1}$, and $25$ gradient steps. The EMAIL dataset configuration used similar layer targeting but modified the hyperparameters, with a loss break of $6\times 10^{-4}$, loss prediction probability coefficient of $6.733\times 10^{-1}$, learning rate of $1.868\times 10^{-1}$, and $5$ gradient steps. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ C.6 Max-Entropy Baseline
For the Llama-3-8B model, we targeted layers 33-55 and Patil layers 1515-$31$. The SSN dataset configuration employed a loss break of $1.99\times 10^{-2}$, loss prediction probability coefficient of $20.5$, learning rate of $7.6\times 10^{-3}$, and $25$ gradient steps. The EMAIL dataset adjusted these parameters, using a loss break of $1.75\times 10^{-3}$, loss prediction probability coefficient of $1.35\times 10^{-2}$, learning rate of $9.2\times 10^{-3}$, and $5$ gradient steps. For the GPT-J-6B model, we targeted layers 33-88 and Patil layers 1515-$27$. For the SSN dataset, we used a loss break of $7.0559\times 10^{-2}$, loss prediction probability coefficient of $25.9$, learning rate of $1.937\times 10^{-1}$, and $5$ gradient steps. The EMAIL dataset used a loss break of $3.634\times 10^{-2}$, loss prediction probability coefficient of $1.907\times 10^{-2}$, learning rate of $6.65\times 10^{-2}$, and $10$ gradient steps. In our hyperparameter sweep, we considered the following ranges: learning rate from $5\times 10^{-5}$ to $0.5$, number of gradient steps from 5 to 25, loss break from $1\times 10^{-4}$ to $0.1$, loss prediction probability coefficient from $0.1$ to $100$, and different layer selection.
@@ D.1 Alternative Neuron Selection Strategies
We explored several alternative methods for neuron selection, besides our hybrid approach that considers both neuron activation and association with the target token (Section˜3.1.2). These alternatives include: (a) selecting the most activated neurons given the prompt, (b) selecting neurons solely based on their association with the target token, (c) neurons with highest gradients with respect to the prompt (Dai et al., 2022), and (d) random neuron selection as a baseline. As shown in Table˜7, our evaluation on the Emails dataset on GPT-J-6B, using the same seed (0) that was used for finding the best hyper-parameters, demonstrated that the hybrid approach (rank & activation) outperformed these alternatives in achieving superior results in both metrics of unlearning effectiveness and model integrity.
@@ D.2 Alternative Neuron Editing Strategies
We explore an alternative neuron editing strategy inspired by DEPN, where the target neurons are directly zeroed out instead of applying our proposed rank-based editing approach. As shown in Table˜8, zeroing neurons results in significantly inferior performance compared to REVS.
@@ E.2 Alternative Token Selection Strategies
We investigated token selection strategies for unlearning on the GPT-J-6B Email dataset. Table˜9 shows that selecting the rarest tokens yields the best performance across all metrics. The rarest token approach significantly outperforms other strategies like most frequent, first, or random token selection, demonstrating substantially higher effectiveness in unlearning.
@@ F.2 Synthetic SSN Dataset
We then randomly assigned fake sensitive information to each placeholder. To populate the placeholders, we used Claude 3 Sonnet to generate a list of $40$ unique names (since each sentences will be assigned with the same name and SSN for allowing evaluating generalization). Additionally, we generated $200$ random dates and $40$ random Social Security Numbers (SSNs). To get the model to memorize the sentences containing the sensitive information, we fine-tuned the model with a low learning rate of 2e62e-6 to minimize unintended impact. The fine-tuning process ran for 4 epochs with a batch size of $1$. Table˜12 shows 10 example sentences from this dataset.